Grade 11General Physics

what will be the pressure exerted by the earth 's atmosphere at the sea level of measured in cgs system of unit?

riya
8 Years agoGrade 11

1 Answer

Shailendra Kumar Sharma
8 Years ago
Pressure at sea level is 1 atm = 101325Pa =101325N/m2 =101325 kg/(s2*m) =101325* (1000g)/(s*100cm) =1013250g/cm.s2